OpenStudy (anonymous):

MEDAL+FAN+TESTIMONIAL Enter an equation in standard form to model the linear situation. A barrel of oil was filled at a constant rate of 7.7 gallon/minute. The barrel had 11 gallons before filling began.

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

11 will be your y-intercept and 7.7 will be your slope

OpenStudy (anonymous):

Because it already had 11 and fills at 7.7 gallons/min

OpenStudy (welshfella):

y = mx + b m = slope, b = y intercept

OpenStudy (anonymous):

So substitute into y=mx+b

OpenStudy (tsukiyama_shuu):

So it is Y=7.7X + 11 -7.7X + Y = 11 In standard form.

OpenStudy (whpalmer4):

@Tsukiyama_Shuu Often standard form is taken to mean that the coefficients of x and y are integers, and that the coefficient of x is positive.

OpenStudy (tsukiyama_shuu):

@whpalmer4 So what do you mean? What would I have to change to make my statement correct?

OpenStudy (anonymous):

That is slope intercept form not standard

OpenStudy (anonymous):

Just change the terminology

OpenStudy (whpalmer4):

well, you could multiply the entire equation by -10, which would give: \[(-10)(-7.7x) + (-10)y = (-10)11\]\[77x - 10y = -110\]

OpenStudy (tsukiyama_shuu):

Oh? I'm pretty sure that I was taught that when X and Y are on the same side, its the standard form.. Oh well, I'm probably wrong on that, It's been awhile.

OpenStudy (whpalmer4):

That would be standard form in the most conservative definition.

OpenStudy (tsukiyama_shuu):

Oh I see.

OpenStudy (whpalmer4):

it works exactly the same as you version, of course...just a question of definition

OpenStudy (tsukiyama_shuu):

Aye, true. I suppose that it's a miracle that I even got it right xD

OpenStudy (whpalmer4):

@Brill making a comment like "that is slope intercept form not standard" in the middle of a conversation without making it clear what you are referring to by "that" tends to confuse people...especially those people who don't already know the material well, and even some of those who do :-)